### v1

# FinEval-KR: A Financial Domain Evaluation Framework for Large Language Models’ Knowledge and Reasoning

Abstract
Large Language Models (LLMs) demonstrate significant potential but face challenges in complex financial reasoning tasks requiring both domain knowledge and sophisticated reasoning. Current evaluation benchmarks often fall short by not decoupling these capabilities indicators from single task performance and lack root cause analysis for task failure. To address this, we introduce FinEval-KR, a novel evaluation framework for decoupling and quantifying LLMs’ knowledge and reasoning abilities independently, proposing distinct knowledge score and reasoning score metrics. Inspired by cognitive science, we further propose a cognitive score based on Bloom’s taxonomy to analyze capabilities in reasoning tasks across different cognitive levels. We also release a new open-source Chinese financial reasoning dataset covering 22 subfields to support reproducible research and further advancements in financial reasoning. Our experimental results reveal that LLM reasoning ability and higher-order cognitive ability are the core factors influencing reasoning accuracy. We also specifically find that even top models still face a bottleneck with knowledge application. Furthermore, our analysis shows that specialized financial LLMs generally lag behind the top general large models across multiple metrics.
FinEval-KR: A Financial Domain Evaluation Framework for Large Language Models’ Knowledge and Reasoning
Shaoyu Douthanks: Equal contribution. Ant Group                        Yutian Shenthanks: This work was completed during the internship at Ant Group. , Mofan Chen∗† Zixuan Wangand Jiajie Xu Shanghai University of Finance and Economics
Qi Guo, Kailai Shao Chao Chen  and Haixiang Hu Ant Group                        Haibo Shi Min Min  and Liwen Zhangthanks: Corresponding author. Shanghai University of Finance and Economics
1 Introduction
In recent years, rapid LLM development has led the transformation of artificial intelligence. LLMs demonstrate strong natural language processing capabilities and inspire application innovations across various fields, including scientific research Zhang et al. (2024), financial services Nie et al. (2024b), content creation Betker et al. (2023), etc. However, achieving satisfactory performance for complex tasks, such as financial decision making, proves difficult when relying solely on knowledge introduced during training or instructions Wang and Brorsson (2025); Liu et al. (2025); Wang et al. (2024). Reasoning ability, that is, the ability of logical deduction, problem solving, and abstract thought, stands as a core hallmark of advanced intelligence and becomes crucial for assessing LLM intelligibility and applicability Wang and Song (2024); Li et al. (2024c); Valmeekam et al. (2023b). Therefore, it is critical to accurately assess the knowledge and reasoning abilities of LLMs. This helps to understand the shortcomings of the model to support targeted optimization.
Although several LLM evaluation benchmarks have been proposed, they still have several limitations in evaluating reasoning capabilities.
Insufficient capability decoupling. Mainstream benchmarks typically evaluate a model’s capabilities based on its performance across various tasks. Their performance represents either the model’s knowledge capacity Nie et al. (2024a); Liu and Jin (2024) or its reasoning ability Saparov and He (2023); Geva et al. (2021). However, our experiments demonstrate that LLM performance in reasoning tasks is influenced by both knowledge and reasoning ability. Therefore, it is essential to decouple and quantify them separately to achieve a more accurate capability characterization.
Lack of root cause analysis. Current reasoning evaluation frameworks mostly focus on the correctness of the reasoning processes and results, while neglecting the diagnosis of the erroneous result. Specifically, there is not yet an effective way to distinguish whether a reasoning failure stems from knowledge gaps, such as unclear concept comprehension, or from flaws in the reasoning processes, like missing reasoning steps. This limits the potential for specifically optimizing the model.
Neglect of the cognitive science perspective in financial LLM benchmark. Turning the perspective to the financial domain, a majority of work concentrates on assessing model performance in financial NLP tasks and knowledge coverage, with only a few focusing the reasoning ability. Even so, these methods lack systematic design from a cognitive science perspective, making it challenging to uncover differences in model capabilities across various cognitive levels.
The above limitations motivate us to explore the following questions:
Q1: How do knowledge and reasoning ability jointly determine the performance of LLMs in domain reasoning tasks?
Q2: Can we develop an evaluation framework that decouples and independently quantifies knowledge and reasoning ability from task performance?
To address these questions, this paper first conducts a preliminary experiment taking the financial domain as an entry point to empirically verify the fundamental role of knowledge in domain reasoning tasks (see Appendix A for details). Furthermore, to bridge the gap in relevant research and provide reproducible research resources, the paper makes the following main contributions:
  • A novel evaluation framework. We design a novel LLM evaluation framework that, for the first time, effectively separates domain knowledge assessment and reasoning assessment from a sole task performance metric, and proposes novel metrics: knowledge score and reasoning score. Moreover, inspired by cognitive science, a prerequisite for an LLM to solve complex reasoning problems is that it possesses a certain level of “cognitive abilities”. This motivates us to propose another new metric, cognitive score, based on Bloom’s cognitive taxonomy, to explore in a fine-grained manner the abilities exhibited by LLMs when facing reasoning tasks.
  • A new open-sourced financial domain reasoning dataset [^1] The dataset will be open-sourced after acceptance of the paper.. The released dataset encompasses 22 key financial subfields and provides sample-level knowledge labels, step-level reasoning processes, and cognitive abilities required for the tasks. The dataset contributes to the community by providing a comprehensive and specialized benchmark for research in financial and cognitive reasoning.
  • Evaluation results of mainstream LLMs. Based on the proposed framework and dataset, we conduct a comprehensive evaluation of the current mainstream LLMs, which verifies the effectiveness of the proposed evaluation methodology and yields a series of insightful conclusions (see Section 4.3 and Appendix G for details).
2 Related Work
2.1 Current LLM Benchmarks
2.1.1 Evolution of LLM benchmarks
Recent studies have identified a paradigmatic shift in LLM benchmarks, transitioning from task-specific evaluation to capability-based evaluation that measure core competencies including knowledge, reasoning, safety, etc. Cao et al. (2025). It is also worth noting that many works combine the two approaches. This section first summarizes the two evolutionary phases to clarify the developmental trajectory, and then focuses on reviewing knowledge and reasoning evaluation methods that are particularly relevant to this work.
Task-specific Evaluation. Early LLM benchmarks extended traditional NLP evaluation paradigms to establish baseline performance metrics. The GLUE benchmark Wang et al. (2019b) and its successors, e.g., SuperGLUE Wang et al. (2019a), CLUE Xu et al. (2020), FewCLUE Xu et al. (2021), developed a set of core NLU tasks encompassing sentiment analysis, similarity analysis, etc. While these benchmarks provided crucial standards for early LLM development, their scope proved insufficient to capture advanced capabilities such as reasoning ability, common-sense comprehension with the evolution of LLMs.
Multidisciplinary Knowledge Evaluation. The MMLU benchmark Hendrycks et al. (2021) represented a paradigm shift by systematically evaluating cross-domain knowledge across 57 distinct tasks, spanning STEM disciplines, humanities, and social sciences. This multidisciplinary approach was further adapted to multilingual contexts through works like C-Eval Huang et al. (2023). Similar studies include those by Hu et al., 2024; Yu et al., 2023. Simultaneously, domain-specific evaluation frameworks emerged to address specialized knowledge requirements, particularly in fields such as medicine Pal et al. (2022); Shah et al. (2025) and law Li et al. (2024a); Fei et al. (2024), as they recognized the limitations of generic assessments in capturing discipline-specific expertise.
Reasoning Evaluation. Recently, LLM benchmarks have started to prioritize the evaluation of reasoning capabilities. The Chain-of-Thought Hub Fu et al. (2023) pioneered systematic assessment of multi-step reasoning across mathematical, symbolic, and knowledge-intensive domains, establishing reasoning capacity as a critical differentiator between weaker and stronger LLMs. PlanBench Valmeekam et al. (2023a) was developed to assess LLMs’ planning and reasoning capabilities about change, providing sufficient diversity in both task domains and planning capabilities. Similar studies include those by Gui et al., 2024; Wang et al., 2023a; Zhou et al., 2024; Sprague et al., 2023; Shi et al., 2023; Cao et al., 2024 Furthermore, the LLM-agent-focused evaluation framework is now at the forefront of LLM evaluation Shen et al. (2024); Abdelnabi et al. (2024); Madge and Poesio (2024).
2.1.2 Paradigms and Limitations in Evaluating LLM Reasoning Capabilities
Current approaches to assess LLMs’ reasoning capabilities exhibit the following two paradigms.
Outcome-Oriented Metrics. The most straightforward methodology relies on task-specific accuracy metrics. Mathematical reasoning benchmarks like GSM8K Cobbe et al. (2021) have become particularly influential. Although this method provides direct, quantifiable indicators, it closely combine model knowledge and reasoning, making it unclear which aspect is being measured. We also confirmed this statement through a quantitative experiment, as detailed in Appendix A. This led us to develop a new evaluation framework and metrics that can decouple knowledge and reasoning from task accuracy.
Process-Oriented Metrics. A risk of the previous approach is the Clever Hans effect Niven and Kao (2019). To address this, researchers have explored methods to transparentize the reasoning process. The ProofWriter dataset Tafjord et al. (2021), which contains small rulebases of facts and rules along with questions that can be proven true or false, enables evaluation of models’ intermediate reasoning steps. Similarly, code-based reasoning evaluations through Python execution reveal the logical progression of problem-solving Chen et al. (2023); Yang et al. (2025). These transparent reasoning paths also build the foundation for root cause analysis of reasoning errors. These works collectively demonstrate that step-level verification is necessary for reasoning evaluation.
2.2 LLM benchmarks in Finance
Financial LLM benchmarks have progressed through two development phases mirroring general LLM benchmark evolution while addressing domain-specific requirements. Initial financial benchmarks adapted general NLP tasks to financial contexts, representative work includes: FinGPT Wang et al. (2023b), CFBenchmark Lei et al. (2023), CFLUE Zhu et al. (2024) and Golden-Touchstone Wu et al. (2024b). As the complexity of the tasks in this domain gradually increased, the key focus of the evaluation methods evolved to test specialized knowledge competencies, such as FinEval Zhang et al. (2023), UCFE Yang et al. (2024), MTFinEval Liu and Jin (2024), CFinBench Nie et al. (2024a) and FintruthQA Xu et al. (2024). Contemporary benchmarks, e.g., MME-Finance Gan et al. (2024), InvestorBench Li et al. (2024b), and FinBen Xie et al. (2024) now incorporate decision-making tasks that require integrative reasoning in financial scenarios, which involve market analysis, risk assessment, and investment strategy development.
Financial reasoning tasks differ significantly from general reasoning tasks in the following three aspects: (1) complex data comprehension, e.g., numerical time series like stock prices and unstructured text from earnings reports; (2) domain-specific knowledge, e.g., option pricing formulas and financial regulations; (3) advanced reasoning skills, e.g., formula derivation and numerical computation. These features create a critical limitation in conventional evaluation approaches: that is, overall performance metrics often fail to distinguish between knowledge gaps and reasoning deficiencies. This fundamental issue motivates the need for decoupled capability assessment frameworks. Moreover, these differences directly inspired the design of our financial reasoning task dataset, which we designed as a computational problem in a realistic financial scenario (see Section 3.1 for details).
Existing work suggests that evaluating LLMs from a cognitive science perspective, which is more fine-grained than the task-based approaches, often offers crucial insights for LLM optimization Huber and Niklaus (2025); Adams (2015). However, most financial LLM benchmarks currently overlook such fine-grained analysis. This gap becomes particularly problematic when engineers want to fine-tune a base model to make targeted improvements, as training dataset is constructed based on a fine-grained measure of the base model’s capabilities. In summary, a more fine-grained evaluation and diagnostic approach helps to identify specific development needs in the model training pipeline.
3 FinEval-KR
In this section, we present a Financial domain Evaluation framework for assessing Knowledge and Reasoning abilities (FinEval-KR). We first describe the methodology used to construct the evaluation dataset. Next, we introduce a multi-stage evaluation framework that performs root cause analysis of reasoning errors via knowledge-augmented question answering, enabling a decoupled assessment of a model’s knowledge mastery and reasoning ability. Finally, we define a series of evaluation metrics: first, a knowledge score based on domain knowledge coverage; second, a reasoning score; and third, a cognitive score which is based on Bloom’s taxonomy. This approach improves the interpretability of LLM evaluations in financial scenarios and provides clear directions for targeted model improvement.
3.1 Benchmark Dataset Construction
The FinEval-KR benchmark dataset is constructed through four steps: data collection and processing, automated question generation, answer generation, and dataset annotation. This framework ensures comprehensive coverage of financial knowledge domains while maintaining academic rigor. All prompt templates for the dataset generation are detailed in Appendix B.
Data Collection and Processing
We select nine canonical textbooks across major financial disciplines. These disciplines together provide a comprehensive and up-to-date view of finance, and thus ensure the dataset is both authoritative and relevant. Please refer to Appendix C for detailed data sources and reasons for selection.
The dataset construction process follows a three-stage workflow. First, content extraction employs OCR technology to identify and convert textual content and mathematical equations from source materials into machine-readable Markdown format. Subsequently, the content cleaning phase involves manual removal of non-essential components (e.g., prefaces, appendices) combined with quality assurance checks to eliminate irrelevant information. Finally, standardization transforms the cleaned data into structured data. This sequential pipeline ensures data integrity while maintaining compatibility with automated processing requirements.
Automated Question Generation
The automated question generation process consists of the following steps. First, the corpus construction phase constructs 8,460 pages of financial content covering all nine textbooks. Second, the custom-designed prompt ensures that the generated questions meet specific criteria (see Figure 6 in the appendix for details). Third, the multi-stage validation phase utilizes GPT-o1 to perform checks such as logical validation, consistency assessment, and final validation, whereby only questions that pass all three checks are kept in the final dataset.
Answer Generation
The answer generation process follows a three-step approach. Step 1: answers are initially generated without format constraints to capture diverse problem-solving approaches. Step 2: these raw solutions then undergo standardized formatting according to predefined templates to ensure uniform presentation. Step 3: a triple-validation protocol, mirroring the question validation, is applied, with GPT-o1 adopted at this stage..
Dataset Annotation
Then we use GPT-o1 with prompt engineering to annotate knowledge points involved in each question, and corresponding Bloom’s taxonomy levels for each solution step. Knowledge annotations include, but are not limited to, core financial concepts, regulatory frameworks and mathematical operations. For Bloom’s taxonomy annotation, we map each step to one or more of the following levels: remembering, understanding, applying, analyzing and evaluating. Note that the creation is excluded because all questions in this dataset are objective finance problems with fixed answers that cannot measure creativity. Additionally, creative responses would introduce subjective bias during assessment, and this aligns with design principles of current standardized tests like the qualification certificate of banking professional, licensing examination for fund practitioners, etc.
Dataset Characteristics and Exemplary Samples
The final evaluation set contains a total of 9,782 question-answer pairs and their associated annotations. The questions and answers in the dataset are verified by graduate students with backgrounds in finance, economics, and statistics, and a sample check showed that the accuracy of the dataset is above 90%. An exemplary sample from the constructed dataset is given in Figure 1 on a yellow background. (see Figure 10 for the complete example). The complete statistical characterization of the dataset is detailed in Appendix D.
Figure 1: Three-stage evaluation framework of FinEval-KR, and an exemplary sample of the dataset. Note that the original dataset is in Chinese, the figure provides an English translation for readability.
3.2 Evaluation Framework
As shown in Figure 1, the FinEval-KR framework comprises three progressive evaluation stages. Given an evaluation sample, the first stage requires the evaluated model to answer the question. We then employ an LLM-as-a-judge paradigm,in which a judge model assesses the correctness of the reasoning result. If an error is identified, the judge model pinpoints the first erroneous step and extracts relevant knowledge points. We refer to the output of the judge model as the review result. In the second stage, the knowledge points extracted from the first stage are provided as augmented information to the evaluated model. The model then re-answers the question, and its reasoning result is evaluated by the judge model again. The third stage analyzes the root causes of reasoning errors by comparing the evaluation results from the first and second stages. This ultimately allows for the decoupling and quantification of the model’s knowledge reserve and its reasoning ability.
3.2.1 Stage 1: Question Answering
In this stage, the model under evaluation is required to generate the solution process without format constraints. This design prevents cognitive bias that can arise from forcing alignment with a reference answer’s format. For example, if the evaluated model’s reasoning path has a different order but is logically equivalent to the reference answer (e.g., step 2 in the reference answer is equivalent to the evaluated model’s reasoning step 1), a judge model might tend to compare reasoning steps one by one for consistency, potentially leading to incorrect judgments. For details of the prompt, please refer to Figure 15 in Appendix E.
In FinEval-KR, we utilize another model as the judge model to review the evaluated model’s reasoning process by comparing it with a reference answer. The final evaluation result, as depicted in Figure 2, includes pinpointing the first erroneous step in the reasoning process, identifying its cause, noting corresponding knowledge deficiencies, and determining the overall correctness of the reasoning. Specifically, we prompt the judge model by providing a few examples and employing a Chain-of-Thought (CoT) method. For template details, refer to Figure 14 in Appendix E.
Figure 2: Example of review result generated by the judge model (original in Chinese, with English translation).
It is worth noting that if the evaluated model’s final reasoning result is correct, the judge model considers its entire reasoning process to be correct. This is because our evaluation dataset does not utilize any publicly published or Internet-available questions, and each question necessitates more than two reasoning steps. Consequently, we have good reason to believe that a correct result invariably stems from effective reasoning rather than accidental guessing or the Clever Hans effect.
3.2.2 Stage 2: Knowledge-augmented Answering
If the evaluated model makes a reasoning error in Stage 1, FinEval-KR injects the knowledge extracted in the review result into the prompt and requires the evaluated model to re-answer the question. This approach stems from the conclusion that knowledge is a necessary condition for correct reasoning, which is proven in Appendix A. Subsequently, the judge model employs the same evaluation criteria as in Stage 1 to generate the augmented review result. The prompt template used in this stage is detailed in Figure 16 in Appendix E. Note that we carefully considered the impact of prompt structure on evaluation results during the prompt design. Please refer to Appendix E for details.
3.2.3 Stage 3: Error Diagnosis
Wu et al. (2024a) has demonstrated that the likelihood of an LLM adhering to contextually retrieved information is inversely proportional to the model’s confidence in its non-contextual responses (i.e., prior probability). Therefore, by checking whether the model prefers the external knowledge injected in the Stage 2, we can ascertain whether the relevant knowledge within the model is weak or not. Specifically, the judge model compares the review result from Stage 1 with the augmented review result from Stage 2, following the principles below to determine the root cause of reasoning errors:
  • Knowledge Deficiency. If the augmented review result shows that the model reasons correctly in Stage 2, or the erroneous step occurs later than in Stage 1. This indicates that the evaluated model preferred the augmented knowledge in the second stage. This proves that the initial reasoning error is caused by knowledge deficiency.
  • Reasoning Ability Deficiency. If the evaluated model still makes a reasoning error in Stage 2, and the erroneous step is consistent with Stage 1, this indicates that the evaluated model still preferred its internal prior knowledge. This proves that the initial reasoning error is rooted in poor reasoning ability.
The aforementioned attribution framework achieves the decoupling of knowledge and reasoning ability from task performance metrics.
3.2.4 Evaluation Metrics
We propose three core metrics: namely, knowledge score, reasoning score, and cognitive score, while also retaining accuracy to measure the model’s overall task performance.
Knowledge Score (KS).
This metric quantifies the evaluated model’s knowledge coverage in the financial domain.
$\text{KS}=1-\frac{\left|\bigcup_{i=1}^{M}K^{\prime}_{i}\right|}{\left|\bigcup_{i=1}^{N}K_{i}\right|},$ (1)
where $M$ is the number of errors attributed to knowledge deficiency during evaluation. $K^{\prime}_{i}$ denotes the set of knowledge points involved in the erroneous reasoning steps for the $i$-th question whose error was attributed to a knowledge deficiency. $N$ is the total number of evaluation samples. $K_{i}$ denotes the set of knowledge points involved in the $i$-th evaluation question. The denominator in the Eq. (1) is the total number of knowledge points across all evaluation questions in the dataset.
Reasoning Score (RS).
This metric measures the evaluated model’s reasoning ability in the financial domain.
$\text{RS}=\frac{\sum_{i=1}^{N}\mathbb{I}(a_{i}=a_{i}^{\text{ref}})}{N-\sum_{i=1}^{N}\mathbb{I}(a_{i}\neq a_{i}^{\text{ref}}\land r(a_{i})=\text{knowledge})},$ (2)
where $a_{i}$ is the answer of the evaluated model for the $i$-th question, and $a_{i}^{\text{ref}}$ is the corresponding reference answer. Thus, the numerator of Eq. (2) is the total number of questions with correct reasoning. $r(a_{i})$ represents the root cause for the erroneous reasoning result $a_{i}$. The output of $r(\cdot)$ can be knowledge or $\text{reasoning}^{j}$, $j\in\{1,2,3,4,5\}$, where $j$ is the index of the level in Bloom’s taxonomy, and $1,2,3,4,5$ denote remembering, understanding, applying, analyzing, and evaluating respectively. The denominator in Eq. (2) is the total number of questions whose reasoning errors are not attributed to knowledge deficiencies.
Cognitive Score (CS).
Furthermore, to explore the level of cognition exhibited by LLMs when solving reasoning tasks, we expanded a series of fine-grained metrics drawing on RS, based on the assumption that LLMs need to have a certain level of cognition to conduct correct reasoning. The $j$-th cognitive level score of the LLM is defined as,
$\text{CS}_{j}=RS\times(1-$ (3)
$\alpha\times\frac{\sum_{i=1}^{N}\mathbb{I}(r(a_{i})=\text{reasoning}^{j})}{N-\sum_{i=1}^{N}\mathbb{I}(a_{i}\neq a_{i}^{\text{ref}}\land r(a_{i})=\text{knowledge})}),$
where $j\in\{1,2,3,4,5\}$ and $\alpha\in(0,1)$ is a penalty coefficient, inversely proportional to the index $j$; this means that the absence of lower-level reasoning abilities leads to more severe deductions. The reason RS is used as a multiplier in Eq. (LABEL:eqn:rs_j) is that all cognitive deficits lead to reasoning failures.
Accuracy (Acc).
Finally, the success rate for all reasoning tasks is defined as,
$\text{Acc}=\frac{\sum_{i=1}^{N}\mathbb{I}(a_{i}=a_{i}^{\text{ref}})}{N}.$ (4)
4 Experiments
This section first validates the FinEval-KR framework’s effectiveness in identifying potential knowledge weaknesses in evaluated models and performing root cause analysis through an empirical study. Subsequently, we selected a range of LLMs, both open-source and closed-source, along with some LLMs with small parameters, and evaluated them using the proposed FinEval-KR framework and our open-source dataset.
4.1 Alignment Analysis of Judge Model in FinEval-KR
The core of FinEval-KR lies in the design of the judge model for identifying potential knowledge weaknesses and pinpointing root causes, which is fundamental for achieving decoupled evaluation. To this end, we design a set of comparative experiments to assess the alignment between the judge model and human evaluation.
The methods participating in the comparison include: (a) Direct Prompting (Direct): we use advanced models like GPT-o1, directly prompting them to execute knowledge identification and root cause localization tasks. (b) Task Decomposition (Decomp.): decompose the above two tasks into multiple subtasks and clearly define the logical dependencies between these subtasks (c) Task Decomposition with CoT (Decomp. with CoT): based on task decomposition, this method requires the model to self-reflect and explicitly output reasoning processes step-by-step before generating the final conclusion (i.e., adding the <Inner Thought> as shown in Figure 14).
Note that the latter two methods are implemented using Qwen2.5-72B_Instruct (see Figure 14 in the appendix for details). We use the following three metrics to evaluate the performance of these methods: (a) Localization Accuracy ($\text{Acc}_{\text{loc}}$): this refers to the proportion of correctly localized first incorrect step in the reasoning process. (b) Recall of Knowledge Point (Recall): the percentage of missing knowledge correctly identified in Stage 1. (c) Attribution Accuracy ($\text{Acc}_{\text{attr}}$): The proportion of correctly attributed reasoning errors to knowledge or reasoning ability deficits.
All metrics are calculated after a manual review of the model outputs by financial domain experts. Experimental results are presented in Table 1. It can be observed that FinEval-KR outperforms the comparison methods across all metrics, demonstrating its superior performance in error attribution and knowledge identification tasks.
Meanwhile, the experimental results show that FinEval-KR’s automated evaluation method has a certain degree of error (i.e. about 8%, 6% and 7% errors in localization, knowledge retrieval, and attribution, respectively), but we believe that this level of error is acceptable and the evaluation results are discriminative. In addition, we discuss the bias challenges encountered by the judge model in the experiments and their countermeasures in Appendix F.
Methods/Metrics $\text{Acc}_{\text{loc}}$ Recall $\text{Acc}_{\text{attr}}$
Direct 0.56 0.24 0.20
Decomp. 0.85 0.60 0.53
Decomp. with CoT (ours) 0.92 0.94 0.93
Table 1: Performance metrics for three root cause analysis methods, where higher numbers mean better performance.
4.2 Evaluation
To comprehensively evaluate the knowledge and reasoning capabilities of LLMs in the financial domain, we selected 23 representative LLMs that are currently highly popular, by referencing Chatbot Arena [^2] https://openlm.ai/chatbot-arena/. These models cover a variety of types, including open-source and closed-source, different parameter scales, and various architectures such as dense and MoE. The specific version of the selected model is detailed in Appendix F.
4.3 Results and Core Findings
Table 2 lists the performance of the 23 LLMs on the FinEval-KR dataset. Results in the table are from a single run of each model. Since metrics $\text{CS}_{1}$ and $\text{CS}_{2}$ do not show significant variation among models, they are omitted here. The complete results can be found in Table 5 in Appendix G.
Model/Metrics Acc KS RS
$\text{CS}_{3}$
$\text{CS}_{4}$
$\text{CS}_{5}$
Open-source lightweight LLMs without reasoning
GLM-4-9B 0.284 0.782 0.396 0.072 0.396 0.393
Mistral-small-22B 0.378 0.814 0.509 0.168 0.509 0.505
Qwen2.5-14B_Instruct 0.548 0.849 0.688 0.394 0.688 0.683
Open-source lightweight LLMs with reasoning
Qwen2.5-14B-R1-distilled 0.640 0.864 0.785 0.555 0.785 0.780
QwQ-32B-preview 0.731 0.901 0.850 0.674 0.850 0.847
Open-source LLMs without reasoning
Qwen2.5-72B_Instruct 0.641 0.886 0.769 0.525 0.769 0.749
DeepSeek-V3 (671B) 0.740 0.925 0.835 0.654 0.811 0.833
Open-source LLMs with reasoning
DeepSeek-R1 (671B) 0.860 0.960 0.928 0.840 0.917 0.926
Closed-source LLMs without reasoning
Doubao-pro-32k 0.424 0.805 0.557 0.229 0.557 0.553
Moonshot-v1-8k 0.479 0.840 0.613 0.289 0.612 0.610
Moonshot-v1-128k 0.453 0.837 0.593 0.269 0.557 0.589
Ernie-Bot-4.0 0.571 0.872 0.697 0.422 0.693 0.664
Qwen-max-latest 0.647 0.885 0.770 0.534 0.742 0.767
GPT-3.5-turbo 0.287 0.757 0.398 0.093 0.398 0.394
GPT-4o 0.669 0.884 0.799 0.578 0.798 0.795
GPT-4.1 0.824 0.952 0.905 0.788 0.891 0.904
Closed-source LLMs with reasoning
GPT-o1-mini 0.751 0.903 0.847 0.703 0.847 0.845
GPT-o3-mini 0.810 0.914 0.901 0.800 0.890 0.899
Gemini-2.5-pro 0.884 0.978 0.896 0.807 0.875 0.884
Gemini-2.5-flash 0.840 0.956 0.910 0.824 0.898 0.905
Claude-3.7-sonnet 0.788 0.939 0.873 0.731 0.855 0.871
Financial LLMs with reasoning
Xuanyuan-FinX1-preview 0.587 0.865 0.731 0.460 0.731 0.728
Fin-R1-7B 0.412 0.731 0.553 0.214 0.553 0.548
Table 2: Reasoning Accuracy (Acc), Knowledge Score (KS), Reasoning Score (RS), and Cognitive Score (CS) of evaluated LLMs on the FinEval-KR. The complete results please refer to Table 5 in the appendix.
From the results in Table 2, we draw the following core findings. The analysis of the model performance echelons and discussion of the results are detailed in Appendix G.
Common Characteristics of Tier 1 Models
A comprehensive analysis of all evaluation metrics identifies the current Tier 1 models as Gemini 2.5 Pro/Flash, DeepSeek-R1 and GPT-4.1. These models typically have parameters in excess of a hundred billion and use the MoE architecture to optimize computational resources. Furthermore, they specifically optimize reasoning capabilities through methods like reinforcement learning, and demonstrate outstanding performance in knowledge coverage and the completeness of reasoning paths.
Relationships Between Evaluation Metrics
KS shows a significant positive correlation with RS. However, among top models, KS differences tend to converge. This indicates that knowledge coverage is no longer the performance bottleneck for mainstream LLMs. RS has a strong correlation with accuracy. This demonstrates that a model’s reasoning ability is crucial for successful reasoning. From a cognitive science perspective, the ability to apply knowledge ($\text{CS}_{3}$) serves as a dividing line for model performance differentiation. $\text{CS}_{3}$ drops sharply in Tier 3 models, which directly impacts overall RS and accuracy. In summary, KS measures basic knowledge ability, while RS and higher-order cognitive metrics ($\text{CS}_{3}$ to $\text{CS}_{5}$) are the core driving factors for improving accuracy.
Bottleneck in Higher-Order Cognitive Abilities
Even for top models, their $\text{CS}_{3}$ scores remain significantly lower than their $\text{CS}_{4}$ and $\text{CS}_{5}$ scores. For example, GPT-4.1 has a $\text{CS}_{3}$ score of 0.824 compared to a $\text{CS}_{4}$ score of 0.910. This demonstrates that existing LLMs still have clear shortcomings in knowledge transfer and practical real-world application.
Development Direction for Financial LLMs
Xuanyuan-FinX1-preview consistently outperformed Fin-R1-7B, primarily due to its larger parameter size (72B). It also demonstrated a clear advantage over models released before Q4 2024 (when its base model, Qwen2.5-72B_Instruct was released). However, Xuanyuan-FinX1-preview currently lags behind advanced general LLMs by approximately 25% in financial reasoning tasks. This gap highlights limitations in data breadth, domain generalization, and iteration efficiency for industry-specific models, indicating significant future optimization potential.
5 Conclusion
This paper introduces FinEval-KR, a novel evaluation framework designed to decouple and assess the knowledge and reasoning abilities of LLMs in the financial domain, supplemented by a cognitive perspective and a new dataset. Our evaluation results indicate that reasoning and higher-order cognitive abilities are crucial for reasoning accuracy. Even top models encounter a bottleneck in knowledge application, and specialized financial models generally lag behind top general LLMs.
Limitations
In this study, we currently use Qwen2.5-72B_Instruct as our judge model based on the following three considerations.
  1. 1.
    Knowledge Coverage and Semantic Understanding. As one of the largest open-source Chinese LLMs, Qwen2.5-72B_Instruct possesses an extensive financial knowledge base. It demonstrates excellent performance in terminology accuracy and cultural context adaptability.
  2. 2.
    Evaluation Stability and Consistency. Preliminary tests indicate that Qwen2.5-72B_Instruct, when serving as the judge, strictly adheres to preset standards. Its output maintains high consistency across repeated experiments and shows strong concordance with human evaluations.
  3. 3.
    Deployment Cost and Efficiency. Due to its open-source nature, we can deploy the entire evaluation system on a local system. This significantly increases processing speed and reduces budget consumption while ensuring assessment quality.
While specialized reasoning models like DeepSeek-R1 can empirically increase evaluation accuracy, outputting their reasoning process significantly increases evaluation time. Given this trade-off, we deem the current 92% accuracy (as mentioned in Section 4.1) acceptable and choose not to use other reasoning models.
However, Qwen2.5-72B_Instruct may limit overall evaluation accuracy and quality. Therefore, in the future, we plan to introduce more powerful LLMs (e.g., GPT-o1) or incorporate human-in-the-loop paradigms. These additions further improve the quality of the evaluation results. Consequently, we will release updated benchmarks accordingly. Similarly, the accuracy of the dataset samples can be improved by adding manual calibration and feedback.
Finally, to assess model capabilities, we currently employ the computational problem format in the FinEval-KR benchmark. Note that our framework can easily extend to other benchmark forms like multiple-choice and expository questions, representing another future research direction.
Appendix A Preliminary Experimemt
A.1 Experiment Design
To further illustrate our research motivation, we design a set of preliminary experiments focusing on a simple reasoning task in the financial domain – a financial calculation problem that requires only a single formula. In order to correctly solve such problems, LLMs need to complete the following three sub-tasks: (1) Recall the calculation formula for the variable to be solved: assessing the knowledge remembering. (2) Identify the variable names and their values from the problem statement: assessing the knowledge understanding. (3) Substitute the variable values into the formula and calculate the target variable: assessing reasoning ability. According to Bloom’s Taxonomy, these three sub-tasks correspond to remembering, understanding and applying/analyzing respectively. Obviously, the model can only answer the questions correctly if all these subtasks are done correctly.
We design the following three comparative experiments to evaluate the impact of knowledge on reasoning task:
  • Experiment 1 (E1). The LLM is directly prompted with the original question and asked to complete the entire reasoning process independently.
  • Experiment 2 (E2). Based on experiment 1, inject variables and their values into the prompt.
  • Experiment 3 (E3). Further provide formulas for calculation based on experiment 2.
To ensure the fairness, we add an equal amount of irrelevant information as distractors to the control groups, while providing the key knowledge in the experimental groups. The experiments are first conducted on the qwen2.5-7b, and the generalizability of the findings is verified using GPT-4o. The experimental results are presented in Table 3.
Model/Settings E1 E2 E3
Qwen2.5-7B_Instruct 58.0% 72.4% 85.9%
GPT-4o 64.5% 80.5% 92.5%
Table 3: Cumulative correctness rate of reasoning task in three experimental settings.
A.2 Results Analysis
Experiment 3 to 1 can be regarded as knowledge stripping experiments. As key knowledge is progressively removed, the problem-solving rate decreases significantly, suggesting that the lack of knowledge is often the root cause of reasoning failure in reasoning tasks. This leads to the following conclusion.
In complex domain reasoning tasks, knowledge is necessary for successful reasoning.
On the contrary, Experiments 1 to 3 can be regarded as knowledge enhancement experiments, and the results show that the reasoning success rate of LLM is significantly improved after the introduction of key knowledge in the prompts. However, even knowledge is sufficiently injected into the prompt, GPT-4o still persists with an error rate of about 7.5%, suggesting that reasoning ability may become a performance bottleneck for such tasks. Combined with the previous conclusion, we infer that:
Knowledge is a necessary but not sufficient condition for successful reasoning.
It is noteworthy that both models exhibit significant knowledge dependence in all experimental settings, while the performance gap between them persists. This performance discrepancy suggests that there may be significant differences in the knowledge and reasoning capabilities of different models. Therefore, the decoupled evaluation framework can help us identify more clearly the shortcomings of the models in terms of knowledge and reasoning capabilities.
From a cognitive science perspective, knowledge stripping experiments revealed the damaging effects of lower-order cognitive deficits on higher-order reasoning, as evidenced by a 27.94% and 28% decrease in reasoning accuracy in qwen2.5-7b and GPT-4o, respectively. This change also supports the progressive dependence between cognitive levels. Furthermore, the performance difference between the two models in the same experiment settings suggests that there is a significant gap between them, at least in the remembering layer.
A.3 The Dataset and Prompts Used in Preliminary Experimemts
We manually construct a dataset consisting of 200 samples. Each sample in this dataset includes the question, the financial formula being examined, the mapping between the formula’s variables and the specific numerical values, and the ground truth to the question. Figure 3, 4, and 5 show the prompt templates and example samples used in experiments 1, 2, and 3, respectively. In these figures, the content of the prompt template is shown in blue text, while the test samples are shown in black text. The ground truth for this problem is 0.1024 or 10.24%.
Figure 3: The prompt for experiment 1 and an exemplary sample (original in Chinese, with English translation).
Figure 4: The prompt for experiment 2 and an exemplary sample (original in Chinese, with English translation).
Figure 5: The prompt for experiment 3 and an exemplary sample (original in Chinese, with English translation).
Appendix B The Prompt Templates for the Dataset Generation
Figure  6 shows a prompt template for generating questions for a given subfield based on a given piece of corpus. Figure 7 shows a prompt template that generates an solution to a given question. Figures 8 and 9 show the prompt templates for labeling knowledge points and step-level cognitive abilities, respectively.
Figure 6: Prompt template for generating questions for a given subfield based on a given piece of corpus (original in Chinese, with English translation).
Figure 7: Prompt template that generates an solution to a given question (original in Chinese, with English translation).
Figure 8: Prompt template for labeling knowledge points for a given question (original in Chinese, with English translation).
Figure 9: Prompt template for labeling step-level cognitive labels for a given answer (original in Chinese, with English translation).
Appendix C Data Sources for the FinEval-KR Dataset
When constructing a benchmark dataset to evaluate knowledge and reasoning abilities in finance, ensuring the authoritativeness and comprehensiveness of the data sources is paramount. To achieve this, we selected nine widely recognized and classic textbooks from various branches of modern finance. These cover subfield including corporate finance, investments, financial institutions and markets, money and banking, international finance, financial risk management, fixed income securities, financial engineering, and monetary theory and policy. Our goal is to provide a solid theoretical foundation and a wealth of practical examples for the benchmark dataset.
Corporate Finance
  • Selected Textbook: Corporate Finance (13th edition, 2021) by Stephen A. Ross, Randolph W. Westerfield, Jeffrey Jaffe, and Bradford D. Jordan.
  • Rationale: This textbook is widely used in MBA and undergraduate finance courses. It systematically explains core concepts of modern corporate finance, such as arbitrage theory, net present value (NPV), the efficient market hypothesis, agency theory, and the risk-return tradeoff.
  • Covered Financial Subfields: Corporate financing, capital structure, investment decisions, dividend policy, firm valuation, etc.
  • Role in the Benchmark Dataset: Provides a solid theoretical foundation and abundant practical examples for reasoning and computational questions in corporate finance.
Investments
  • Selected Textbook: Investments (13th edition, 2023) by Zvi Bodie, Alex Kane, and Alan J. Marcus.
  • Rationale: This book deeply explores securities market efficiency, risk-return relationships, and asset allocation strategies. Its content is highly aligned with the CFA (Chartered Financial Analyst) exam syllabus.
  • Covered Financial Subfields: Securities markets, asset pricing, portfolio theory, behavioral finance, derivatives, etc.
  • Role in the Benchmark Dataset: Offers authoritative theoretical support and practical guidance for reasoning and computational questions in investments.
Financial Institutions and Markets
  • Selected Textbook: Financial Markets & Institutions (13th edition, 2020) by Jeff Madura.
  • Rationale: This book comprehensively analyzes the operational mechanisms and regulatory frameworks of financial institutions like commercial and investment banks. It also provides empirical and case analyses on contemporary hot topics such as stock valuation and market microstructure.
  • Covered Financial Subfields: Financial institutions, financial markets, central banking, monetary policy, market regulation, etc.
  • Role in the Benchmark Dataset: Supplies a systematic theoretical framework and practical examples for reasoning and computational questions concerning financial institutions and markets.
Money and Banking
  • Selected Textbook: The Economics of Money, Banking, and Financial Markets (13th edition, 2021) by Frederic S. Mishkin.
  • Rationale: This book offers an in-depth analysis, from both theoretical and empirical perspectives, of money demand and supply, commercial banking operations and regulation, and the interaction mechanisms between monetary policy tools and financial markets.
  • Covered Financial Subfields: Monetary theory, banking systems, monetary policy, financial markets, etc.
  • Role in the Benchmark Dataset: Delivers in-depth theoretical analysis and empirical support for reasoning and computational questions in money and banking.
International Finance
  • Selected Textbook: International Financial Management (6th edition, 2023) by Jeff Madura and Roland Fox.
  • Rationale: In the context of globalization, this textbook discusses cross-border capital flows, exchange rate volatility, and risk management strategies. It uses numerous case studies to examine practical operations in the international financial environment.
  • Covered Financial Subfields: International capital flows, exchange rate theory, foreign exchange markets, international investment, multinational corporate financial management, etc.
  • Role in the Benchmark Dataset: Provides a global perspective and practical examples for reasoning and computational questions in international finance.
Financial Risk Management
  • Selected Textbook: Risk Management and Financial Institutions (6th edition, 2022) by John C. Hull.
  • Rationale: This book comprehensively reviews methods for measuring and hedging market risk, credit risk, and operational risk. It places particular emphasis on the application of financial derivatives in risk management.
  • Covered Financial Subfields: Risk management, financial derivatives, financial institution regulation, risk measurement and hedging, etc.
  • Role in the Benchmark Dataset: Offers a systematic risk analysis framework and practical guidance for reasoning and computational questions in financial risk management.
Fixed Income Securities
  • Selected Textbook: Fixed Income Securities: Tools for Today’s Markets (4th edition, 2022) by Bruce Tuckman and Angel Serrat.
  • Rationale: This book provides detailed discussions on the pricing principles and trading strategies for fixed income products such as government bonds, interest rate swaps, and credit default swaps.
  • Covered Financial Subfields: Fixed income securities, bond pricing, interest rate derivatives, credit risk, etc.
  • Role in the Benchmark Dataset: Provides authoritative pricing models and practical examples for reasoning and computational questions related to fixed income securities.
Financial Engineering and Derivatives
  • Selected Textbook: Options, Futures, and Other Derivatives (10th edition, 2018) by John C. Hull and Basu Sankarshan.
  • Rationale: This textbook comprehensively covers core topics in financial engineering, including option pricing models, futures contract structures, and the pricing of interest rate and credit derivatives.
  • Covered Financial Subfields: Derivatives markets, option pricing, futures contracts, interest rate derivatives, credit derivatives, etc.
  • Role in the Benchmark Dataset: Offers in-depth theoretical analysis and practical guidance for reasoning and computational questions in financial engineering and derivatives.
Monetary Theory and Policy
  • Selected Textbook: Monetary Theory and Policy (4th edition, 2017) by Carl E. Walsh.
  • Rationale: This book systematically explains the framework of modern monetary theory, focusing on the transmission mechanisms of various monetary policy tools and their effectiveness in low-interest-rate environments.
  • Covered Financial Subfields: Monetary theory, monetary policy, macroeconomic models, policy transmission mechanisms, etc.
  • Role in the Benchmark Dataset: Provides a macroeconomic perspective and policy analysis framework for reasoning and computational questions in monetary theory and policy.
In summary, these nine textbooks are not only authoritative and reliable but also closely aligned with current academic frontiers. They lay a comprehensive and in-depth academic foundation for the financial reasoning and computation benchmark dataset constructed in this study. This ensures that the test questions possess both professional depth and practical relevance.
Appendix D Statistical Characterization of the FinEval-KR Dataset
A complete sample from the constructed dataset is shown in Figure 10. The sample size and its distribution for each subdiscipline in FinEval-KR Dataset is shown in Figure 11, where the subdiscipline categorization methodology refers to Zhang et al. (2023). The top-50 knowledge points in the dataset are shown in Figure 12. In each subdiscipline, the distribution of cognitive labels is shown in Figure 13.
Figure 10: A complete sample from FinEval-KR dataset (original in Chinese, with English translation).
Figure 11: The number of samples in each subdiscipline in the FinEval-KR dataset and their percentage, and “others” in the pie chart includes: econometrics, public finance, insurance, monetary economics, managerial accounting, intermediate financial accounting, corporate strategy and risk management, auditing, cost accounting, taxation and advanced financial accounting.
Figure 12: Top 50 knowledge points in the dataset.
Figure 13: The distribution of cognitive labels in each subdiscipline.
Appendix E Prompt Template Adopted by the FinEval-KR Evaluation Framework
Figure 15 shows the prompt template used in Stage 1 where the evaluated model answers the questions in a free-form format. Figure 16 shows the prompt template used in Stage 2, in which the evaluated model re-answer the question with knowledge point augmented. Note that we have made every effort to ensure the consistency of the prompt structure used in Stage 1 and Stage 2 to avoid changes in reasoning performance due to structural variations. During the prompt design, we conduct multiple rounds of trials and adjustments, ultimately ensuring that changes in the model’s responses are solely related to the injected knowledge points.
Figure 14 shows the prompt templates for the judge model to generate review results and augmented review results based on the reference answers in Stage 1 and 2.
Figure 14: Prompt templates for the judge model (original in Chinese, with English translation).
Figure 15: Prompt template for Stage 1 where the evaluated model answers the questions in a free-form format (original in Chinese, with English translation).
Figure 16: Prompt template for Stage 2 where the evaluated model re-answer the question with knowledge point augmented (original in Chinese, with English translation).
Appendix F Details of Experiment
Bias Challenges in Judge Model
Previous research has shown that using LLMs as judges to evaluate the output of other models inevitably introduces certain evaluation biases, which may lead to unfair comparison results. Therefore, this section will discuss the main evaluation biases discovered during the experimental process of this study and their corresponding mitigation methods.
  • Style Bias: This bias refers to the tendency of the judge LLM to give higher scores to content with a more appealing text style (e.g., clear structure, moderate length), even if the answers contain reasoning errors Wu and Aji (2025). To reduce the impact of this type of bias, we did not restrict the output format of the models being evaluated in Stage 1 and Stage 2, encouraging them to reason freely. Subsequently, we used methods such as regular expressions to unify the original output of each model into a format consistent with the reference answer. This processing method effectively reduced evaluation biases caused by differences in text style.
  • Cognitive Bias: This bias refers to the self-bias that LLMs may exhibit during the evaluation, i.e., a tendency to give higher scores to content they generated themselves, thereby affecting the fairness of the evaluation Koo et al. (2024). To avoid this type of cognitive bias, we excluded GPT-o1 from the scope of evaluated models in our experiment, as it had already been used in the dataset construction and validation. Furthermore, during the preliminary experiment, we tested whether Qwen2.5-72B_Instruct, used as the judging model, exhibited significant cognitive bias in the root cause localization and knowledge gap identification tasks. The experimental results showed that this model did not exhibit a significant self-bias tendency in the aforementioned two tasks. We believe this may be because these two tasks are more specific and objective compared to result comparison tasks without sub-task decomposition, and are further aided by the model introspection prompting shown in Figure 14, which helps enhance the objectivity of Qwen2.5-72B_Instruct during the evaluation.
  • To prevent the judge model from the disturbance of “simple deception” Thakur et al. (2024), we filter out meaningless content in the generated answers, such as isolated affirmative words like “yes” or “of course”, ensuring that the evaluation focuses on the substantive reasoning process rather than superficial linguistic features.
Evaluated Open-source Models
We select several popular LLMs, including DeepSeek-R1, DeepSeek-V3, Qwen2.5-72B_Instruct, and QwQ-32B-preview. To study the effect of model size on performance, we also include smaller models, such as GLM-4-9B, Mistral-small-22B, and Qwen2.5-14B_Instruct. Furthermore, we evaluate Qwen2.5-14B-R1-distilled to investigate the impact of specific reasoning optimization on evaluation metrics. In total, this study includes eight open-source LLMs.
Evaluated Close-source Models
We include a selection of prominent models. This selection covers the latest reasoning models, such as Claude-3.7-sonnet, Gemini-2.5-flash, Gemini-2.5-pro, GPT-o3-mini, and GPT-o1-mini. We also include the current top non-reasoning models: GPT-4.1, GPT-4o, and Qwen-max-latest. Furthermore, we add some models released in 2024 and 2023, including Moonshot-v1-8k, Moonshot-v1-128k, Doubao-pro-32k, Ernie-Bot-4.0, and GPT-3.5-turbo.
Financial LLMs
Furthermore, we specifically select two financial reasoning LLMs for evaluation. The first is XuanYuan-finx1-preview from Duxiaoman AI-Lab, a Chinese financial dialogue and reasoning model designed specifically for the financial domain. It is also the first GPT-o1-like model in the financial industry. The second is Fin-R1, a financial reasoning LLM jointly developed by Shanghai University of Finance and Economics and StepFun Technology. This model is trained on Qwen2.5-7B-Instruct and designed for complex financial reasoning tasks, balancing high performance with low deployment cost.
Implement
During evaluation, all closed-source models are accessed through the official APIs provided by their respective developers. In contrast, open-source models are accessed using the service provided by either Bailian [^3] https://bailian.console.aliyun.comor ModelScope [^4] https://www.modelscope.cn.
Model Version
GLM-4-9B 2024-06-05
Qwen2.5-14B_Instruct 2024-09-19
Qwen2.5-14B-R1-distilled 2025-02-02
QwQ-32B-preview 2025-03-06
Qwen2.5-72B_Instruct 2024-09-19
DeepSeek-V3 (671B) 2025-03-24
DeepSeek-R1 (671B) 2025-01-20
Doubao-pro-32k 2024-06-15
Moonshot-v1-8k 2024-01-31
Moonshot-v1-128k 2024-01-31
Ernie-Bot-4.0 2023-11-17
Qwen-max-latest 2025-01-25
GPT-3.5-turbo 2024-01-25
GPT-4o 2024-11-20
GPT-4.1 2025-04-14
Gemini-2.5-pro 2025-05-06
Claude-3.7-sonnet 2025-02-19
GPT-o1-mini 2024-09-12
GPT-o3-mini 2025-01-31
Gemini-2.5-flash 2025-04-17
Xuanyuan-FinX1-preview 2024-12-27
Fin-R1 2025-03-22
Table 4: Version of the model being evaluated.
Appendix G Discussion of Experimental Results
We evaluate 23 LLMs listed in Appendix F using our proposed financial reasoning dataset. Table 5 presents the complete evaluation results across several metrics: Knowledge Score (KS), Reasoning Score (RS), Cognitive Scores ($\text{CS}_{1}$ to $\text{CS}_{5}$), and Task Accuracy (Acc). The $\text{CS}_{1}$ to $\text{CS}_{5}$ correspond to remembering, understanding, applying, analyzing, and evaluating in Bloom’s taxonomy, respectively.
Model/Metrics Acc KS RS $\text{CS}_{1}$ (remember) $\text{CS}_{2}$ (understand) $\text{CS}_{3}$ (apply) $\text{CS}_{4}$ (analyze) $\text{CS}_{5}$ (evaluate)
Open-source lightweight LLMs without reasoning
GLM-4-9B 0.284 0.782 0.396 0.348 0.356 0.072 0.396 0.393
Mistral-small-22B 0.378 0.814 0.509 0.464 0.474 0.168 0.509 0.505
Qwen2.5-14B_Instruct 0.548 0.849 0.688 0.655 0.659 0.394 0.688 0.683
Open-source lightweight LLMs with reasoning
Qwen2.5-14B-R1-distilled 0.640 0.864 0.785 0.758 0.764 0.555 0.785 0.780
QwQ-32B-preview 0.731 0.901 0.850 0.832 0.838 0.674 0.850 0.847
Open-source LLMs without reasoning
Qwen2.5-72B_Instruct 0.641 0.886 0.769 0.743 0.750 0.525 0.769 0.749
DeepSeek-V3 (671B) 0.740 0.925 0.835 0.816 0.819 0.654 0.811 0.833
Open-source LLMs with reasoning
DeepSeek-R1 (671B) 0.860 0.960 0.928 0.917 0.920 0.840 0.917 0.926
Closed-source LLMs without reasoning
Doubao-pro-32k 0.424 0.805 0.557 0.508 0.515 0.229 0.557 0.553
Moonshot-v1-8k 0.479 0.840 0.613 0.571 0.577 0.289 0.612 0.610
Moonshot-v1-128k 0.453 0.837 0.593 0.550 0.557 0.269 0.557 0.589
Ernie-Bot-4.0 0.571 0.872 0.697 0.661 0.667 0.422 0.693 0.664
Qwen-max-latest 0.647 0.885 0.770 0.746 0.750 0.534 0.742 0.767
GPT-3.5-turbo 0.287 0.757 0.398 0.352 0.361 0.093 0.398 0.394
GPT-4o 0.669 0.884 0.799 0.776 0.780 0.578 0.798 0.795
GPT-4.1 0.824 0.952 0.905 0.894 0.895 0.788 0.891 0.904
Closed-source LLMs with reasoning
GPT-o1-mini 0.751 0.903 0.847 0.833 0.837 0.703 0.847 0.845
GPT-o3-mini 0.810 0.914 0.901 0.893 0.896 0.800 0.890 0.899
Gemini-2.5-pro 0.884 0.978 0.896 0.881 0.883 0.807 0.875 0.884
Gemini-2.5-flash 0.840 0.956 0.910 0.899 0.901 0.824 0.898 0.905
Claude-3.7-sonnet 0.788 0.939 0.873 0.856 0.860 0.731 0.855 0.871
Financial LLMs with reasoning
Xuanyuan-FinX1-preview 0.587 0.865 0.731 0.704 0.711 0.460 0.731 0.728
Fin-R1-7B 0.412 0.731 0.553 0.514 0.521 0.214 0.553 0.548
Table 5: The complete evaluation results across several metrics: Knowledge Score (KS), Reasoning Score (RS), Cognitive Scores ($\text{CS}_{1}$ to $\text{CS}_{5}$), and Task Accuracy (Acc).
G.1 Analysis of Knowledge Score
The KS measures the breadth of a LLM’s knowledge coverage in the financial domain. Based on the evaluation results, models fall into four tiers, as Figure 17 shows.
  • Tier 1 (KS > 0.94) includes primarily top closed-source models (such as Gemini-2.5-pro/flash, GPT-4.1) and the strongest open-source inference model (DeepSeek-R1). These models show extremely high financial knowledge coverage.
  • Tier 2 (KS $\in$ [0.90, 0.94)) includes excellent closed-source models (such as Claude-3.7-sonnet, GPT-o3-mini, GPT-o1-mini) and large-scale open-source models (such as DeepSeek-V3). Their knowledge coverage is slightly lower than Tier 1 but remains at a high level.
  • Tier 3 (KS $\in$ [0.80, 0.90)) covers most mainstream open-source and closed-source models and has a relatively solid foundation in financial knowledge. The specialized financial model Xuanyuan-FinX1-preview is also in this range. This indicates its advantage in financial knowledge acquisition and provides good support for subsequent reasoning tasks.
  • Tier 4 (KS < 0.80) consists mainly of lightweight models (such as GLM-4-9B) or earlier models (such as GPT-3.5-turbo). Their financial knowledge coverage is relatively limited. The specialized financial model Fin-R1-7B is also in this tier. Its lower knowledge score likely results from limitations in parameter size or a lower match with the knowledge required by the test dataset.
In summary, the leading closed-source general large models and some open-source reasoning models perform best in financial knowledge coverage. Model parameter size significantly impacts KS. Current mainstream LLMs are mature in financial knowledge. While knowledge coverage no longer primarily distinguishes top models from other ones, it remains a fundamental ability for high-quality reasoning.
Figure 17: The knowledge score of the models.
G.2 Analysis of Reasoning Score
The RS is inversely related to the proportion of failures caused by incorrect reasoning steps. It reflects a model’s reasoning ability. Based on this metric, the evaluated models are also divided into four tiers, as Figure 18 shows.
  • Tier 1 (RS > 0.90) represents excellent reasoning ability. Except for GPT-4.1, the models in this tier are reasoning-optimized (such as DeepSeek-R1, Gemini-2.5-flash, GPT-o3-mini) and show outstanding performance in reasoning accuracy and the completeness of the logical chain.
  • Tier 2 (RS $\in$ [0.80, 0.90)) shows strong reasoning ability. This tier includes lightweight or earlier reasoning models (such as QwQ-32B-preview, GPT-o1-mini) and some non-reasoning but high-performing models (such as Gemini-2.5-pro, Claude-3.7-sonnet and DeepSeek-V3).
  • Tier 3 (RS $\in$ [0.60, 0.80)) includes models have some reasoning ability but often make more reasoning errors in complex tasks, such as GPT-4o, Qwen2.5-14B-R1-distilled, etc.
  • Tier 4 (RS < 0.60) consists primarily of early non-reasoning models (such as GPT-3.5-turboa, etc.) and lightweight non-reasoning models (such as GLM-4-9B, etc.). These models have a higher error rate during the reasoning.
In summary, explicit reasoning optimization is a key factor that influences the RS and a significant source of current performance differences among models. For example, the reasoning-optimized Qwen2.5-14B-R1-distilled score of 0.785 is significantly higher than the non-reasoning version of 0.688 in the same model family. The results highlight the need to focus future model optimization on the design and optimization of the reasoning pipeline to improve the reliability and stability of executing reasoning tasks.
Figure 18: The reasoning score of the models.
G.3 Analysis of Cognitive Score
The CS provides a systematic evaluation of models’ cognitive abilities based on Bloom’s Taxonomy, across five dimensions, that is remembering ($\text{CS}_{1}$), understanding ($\text{CS}_{2}$), applying ($\text{CS}_{3}$), analyzing ($\text{CS}_{4}$), and evaluating ($\text{CS}_{5}$). As Table 5 shows, CS scores generally exhibit a positive correlation with the KS and the RS.
At the lower cognitive levels ($\text{CS}_{1}$ and $\text{CS}_{2}$), most models achieve high scores, which indicates they have basic remembering and understanding abilities. However, the performance differences widen significantly for higher-order cognitive abilities such as applying, analyzing, and evaluating (i.e. $\text{CS}_{3}$ to $\text{CS}_{5}$). Therefore, our subsequent analysis focuses on these three metrics. We use the average of these higher-order cognitive scores, $\text{CS}_{\text{avg}}$, as the primary evaluation metric. The corresponding performance tiers show in Figure 19.
  • Tier 1 models ($\text{CS}_{\text{avg}}$ >0.85) perform excellently across all cognitive levels and show a distinct advantage particularly in higher-order abilities. DeepSeek-R1 (CS3=0.840, CS4=0.917, CS5=0.926) ranks first and Gemini-2.5-flash (CS3=0.824, CS4=0.898, CS5=0.905) ranks second. Except for GPT-4.1, the models in this tier are reasoning-optimized.
  • Tier 2 models ($\text{CS}_{\text{avg}}\in$ [0.75, 0.85)) also exhibit strong higher-order cognitive abilities, with scores slightly lower than Tier 1. Most excellent general LLMs fall into this range, such as Claude-3.7-sonnet, DeepSeek-V3, etc.
  • Tier 3 models ($\text{CS}_{\text{avg}}\in$ [0.59, 0.75) and CS3 $\in$ [0.4,0.6)) show a significant drop in their $\text{CS}_{3}$ scores compared to Tier 2, and generally fall below 0.60. The financial model Xuanyuan-FinX1-preview is in this tier, which indicates it still has some gap in higher-order cognitive abilities.
  • Tier 4 ($\text{CS}_{\text{avg}}$ < 0.59 and CS3 < 0.4) primarily includes lightweight non-reasoning models, such as GLM-4-9B, with a $\text{CS}_{3}$ of only 0.072. These models are clearly insufficient for higher-order cognitive tasks, where the ability to apply knowledge ($\text{CS}_{3}$) becomes a bottleneck.
In summary, top closed-source models and strong open-source reasoning models demonstrate robust performance across all cognitive levels and show a significant advantage in higher-order cognitive abilities. For most current models, performance in knowledge application ($\text{CS}_{3}$) remains a key bottleneck that limits the improvement of overall cognitive levels.
Figure 19: The average higher-order cognitive scores ($\text{CS}_{3}$ to $\text{CS}_{5}$) for the model.
G.4 Analysis of Task Accuracy
Task Accuracy (Acc) directly reflects the success rate of models in execute reasoning tasks. Only models that possess a broad knowledge base, strong reasoning ability, and good higher-order cognitive skills (especially application and analysis abilities) can achieve high accuracy in practical tasks. The tiers of models based on this metric are shown in Figure 20.
The three best-performing models are Gemini-2.5-pro (0.884), DeepSeek-R1 (0.860), and Gemini-2.5-flash (0.840). These models all exhibit strong reasoning capabilities, and their accuracy is significantly higher than that of non-reasoning models. In contrast, lightweight non-reasoning models such as GLM-4-9B (Acc=0.284) perform the worst. Among the financial LLMs, Xuanyuan-FinX1-preview (0.587) outperforms Fin-R1-7B (0.412), demonstrating the former’s relative advantage in handling domain-specific reasoning tasks.
Figure 20: The accuracy of the models.
G.5 Analysis of the Impact of Model Architecture on Evaluation Metrics
To analyze the impact of model architecture on metrics, we select a set of models using the Mixture-of-Experts (MoE) architecture and models using other architectures such as dense, respectively. Each group includes two open-source and two closed-source models. Their performance lists in Table 6. Note that official documentation on the architectures of Claude-3.7-sonnet and Gemini-2.5-pro is currently unavailable. Based on publicly available information Leanware (2025); OpenAI (2025); Erdil (2024); Howarth (2025); SigmaAI (2025), we categorize them into the Dense or other architecture group for this analysis.
Model/Metrics Acc KS RS $\text{CS}_{1}$ (remember) $\text{CS}_{2}$ (understand) $\text{CS}_{3}$ (apply) $\text{CS}_{4}$ (analyze) $\text{CS}_{5}$ (evaluate)
MoE architecture
DeepSeek-V3 (671B) 0.740 0.925 0.835 0.816 0.819 0.654 0.811 0.833
DeepSeek-R1 (671B) 0.860 0.960 0.928 0.917 0.920 0.840 0.917 0.926
GPT-4.1 0.824 0.952 0.905 0.894 0.895 0.788 0.891 0.904
Gemini-2.5-flash 0.840 0.956 0.910 0.899 0.901 0.824 0.898 0.905
Dense or other architecture
Qwen2.5-72B_Instruct 0.641 0.886 0.769 0.743 0.750 0.525 0.769 0.749
QwQ-32B-preview 0.731 0.901 0.850 0.832 0.838 0.674 0.850 0.847
GPT-o1-mini 0.751 0.903 0.847 0.833 0.837 0.703 0.847 0.845
GPT-o3-mini 0.810 0.914 0.901 0.893 0.896 0.800 0.890 0.899
Table 6: The performance of models for the MoE architecture and Dense or other architectures.
Effects on KS
As shown in Table 6, all of the highest KS (Knowledge Score) results are achieved by models with MoE (Mixture-of-Experts) architectures. While one reason for this might be that MoE models are generally newer, this also confirms the principle that the MoE architecture helps expand a model’s knowledge capacity by activating different expert networks that cover a broader range of knowledge domains. At the same time, compared to Dense architectures, which activate all parameters, MoE only activates a subset of experts during inference, leading to efficient knowledge storage and retrieval.
Effects on RS
In terms of reasoning ability, the MoE architecture is particularly outstanding, it occupies the top three positions. This phenomenon indicates that the MoE architecture appears more suitable for complex reasoning tasks. It suggests that different experts can specialize in specific types of reasoning patterns, and these specialized skills are combined through a dynamic routing mechanism, which enhances the accuracy and robustness of the reasoning chain.
Effects on CS
MoE models show a significant advantage across all cognitive levels, particularly in higher-order abilities such as $\text{CS}_{3}$ to $\text{CS}_{5}$. DeepSeek-R1 and Gemini-2.5-flash nearly dominate the highest scores in these dimensions. This further supports the hypothesis that the MoE architecture handles complex cognitive tasks through expert specialization. May be different types of cognitive operations (e.g., applying knowledge, analyzing reasoning processes, evaluating arguments) are performed by different experts networks.
Effects on Accuray
As a comprehensive metric of knowledge, reasoning, and cognitive abilities, MoE models naturally hold a leading position in task accuracy due to their excellent performance in the aforementioned dimensions. Their multi-expert collaboration mechanism effectively enhances the models’ ability to complete tasks in real-world scenarios.
Appendix H License and Usage Constraints
The released dataset is distributed under the Creative Commons Attribution-NonCommercial 4.0 International License (CC BY-NC 4.0).
The dataset is only for for evaluating LLMs in non-commercial academic research. The dataset is explicitly not authorized for use in training or fine-tuning machine learning models , including pre-training, instruction-tuning, or reinforcement learning stages. Access conditions ensure that all derived data products remain confined to research contexts, with no transfer or application permitted in industrial, governmental, or other operational domains.
Appendix I AI Assistants Usage Disclosure
This study did not employ any AI assistants during the research design, data analysis, or coding phases. During manuscript preparation, the authors exclusively utilized Google Gemini for grammatical refinement and stylistic polishing. No AI-generated content was incorporated into the methodology and results, ensuring the work’s originality and human-driven intellectual integrity.